HUBBLE SPACE TELESCOPE DAILY REPORT #1026 PERIOD COVERED: 0800 EST Daily Status Report as of 335/1200Z. 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WFC 5076 (WF/PC Cycle 3 Calibration: WFC Flats Wide. 2) Routine Earth flat calibration exposures were taken with the Wide Field Camera. The observations were monitored in the OSS, and the data quality was assessed. No problems were reported. 1.2 Completed Two Sets of WFC 4381 (HST Medium - Deep Survey - Cycle 3) The Wide Field Camera was used to make two set of medium-deep high latitude sky survey observations taken in parallel with the prime science observations. The observations were executed as scheduled, and both images were reported to look normal. 1.3 Completed WFC 5269 (Partial Repeat of WF/PC Cycle 2 Interchip Photometric Calibration) The Wide Field Camera was used to observe standard star BD+75D325 to calibrate the DQE (The Detector Quantum Efficiency Factor) of each of the WFC's chips. A total of eight images were received and displayed. All eight images were reported to be normal. 1.4 Completed Two Sets of PC 4770 (WF/PC Cycle 2 Calibration: Internal Monitor Proposal) The Wide Field Camera was used to make bias and dark calibration images. These data are required for routine calibration of the instrument and for data reduction purposes. A total of seven images were received and displayed. All seven of the observations were reported to be normal calibration images. 1.5 Completed PC 4644 (The Nuclei of Nearby S0 and Spiral Galaxies - I A PC Snapshot Imaging) The Planetary Camera was used to make one snapshot observation of the spiral galaxy NGC-7741. The observations used filter F555W and chip PC6. The exposure lasted 300 seconds. The observations was received and displayed, and it was reported to look normal. 1.6 Completed FOS/BL 4673 (Ultraviolet Interstellar Polarization - Cycle 3 High) The Faint Object spectrograph (Blue side) was used to make ultraviolet polarization observations of star HD99872 to make studies of the interstellar medium. The OSS monitored the observations, and assessed the data quality. No anomalies were noted. 1.7 Completed GHRS/2 4541 (Cepheid Masses - Cycle 3 High) The High Resolution Spectrograph was used to make spectral observations of the blue main sequence star which is a companion to Cepheid variable star HD66951. This is to help determine the mass ratio of the two stars. The observations were executed as scheduled, and no problems were reported. 1.8 Completed FOC/96 4674 (" Molecular Emission on HH 47 - Cycle 3 High) The High Resolution Spectrograph was used to make observations of Herbig - Haro Object (HH47). The HRS spectrum should contain a strong florescence feature at 1272 A0 (Driven by Lyman - Alpha) an unidentified line at 1284 A0, and possibly another florescence line at 1293 A0. The observations were executed in Fine Lock with no apparent problems. 1.9 Completed PC 4902 (HST Medium - Deep Survey - Part 2 Low Latitude - Cycle 3 High) The Planetary Camera was used to make a long exposure taken in parallel with the prime science observation. Areas of interest are galactic structure and evolution. There is also the possibility of serendipitous discoveries. The observation was received and displayed in the OSS. It was reported to look okay. 1.10 Completed WFC 4929 (WF/PC Cycle 2: Improved Dark Calibration) The Planetary Camera was used to make two dark calibration images. The two exposures lasted 1200 seconds each, and the camera was pre-flashed before the observations. Both observations were reported to be normal exposures. 1.11 Completed FOC/96 3801 (Spectropolarimetry of Low Red-Shift Active Galactic Nuclei) The Faint Object Camera was used to make observations of quasar 0114-089 to measure the levels of its ultraviolet flux, to see if the flux levels are high enough for subsequent observations using the GHRS and FOS. The observations were executed as scheduled, and no problems were reported. 1.12 Completed FOS/RD 3448 (Ultraviolet Evolution of Elliptical Galaxies at Moderate Redshifts) The Faint Object Spectrograph (Red Side) was used to to make ultraviolet observations of the high red shifted elliptical galaxy A-370-10 to look for changes in the "UV rising branch". This feature is expected to change rapidly with the redshift of the galaxy, and should provide information on the hot stars which are believed to produce it. The observations were monitored in the OSS, and the data quality was assessed. No anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 6 Successful : 6 Scheduled Reacquisitions : 5 Successful : 5 2.2 FHST Updates: Scheduled : 13 Successful : 13 2.3 Significant HSTARs: None 2.4 Operations Note: The launch of STS-61 which was scheduled for this morning Wednesday, Dec.1, 1993 was delayed because of unfavorable weather conditions at the Cape. The new launch window is set for tomorrow, 336/0927Z through 1038Z. Meanwhile, the HST spacecraft continues on with its normal science observations schedule. 3.0 ACTIVITIES SCHEDULED FOR NEXT REPORTING PERIOD: 3.1 Complete Two Sets of PC 5014 (WF/PC Cycle 3 Calibration: PC Flats Narrow. 2) 3.2 Complete PC 5027 (WF/PC Cycle 3 Calibration: Rapid Internal Monitor - Part 2) 3.3 Complete PC 5072 (WF/PC Cycle 3 Calibration: PC Flats Wide. 1) 3.4 Complete Two Sets of 4381 (HST Medium - Deep Survey Cycle 3) 3.5 Complete WFC 5023 (Cal/WFPC Cycle 3 Polarization Calibration) 3.6 Complete GHRS/2 4883 (Weak Absorption Lines in 3C273: Cycle 3 High) 4.0 ANOMALY STATUS: 4.1 Significant New Anomalies: None 4.2 Significant Unresolved Anomalies: None 5.0 SIGNIFICANT FORTHCOMING EVENTS: The launch of STS-61 is scheduled for tomorrow,Thursday, Dec.2, 1993. The new launch window is set for tomorrow, 336/0927Z through 1038Z. /DMH -END-